Ronnie took the first three statistics tests. His scores are the following: 60, 78, and 86. Then, Ronnie took the fourth test. What was his score on the fourth test if his overall average is 79.5?
Answer #1

Mean \bar{x} = \sum X / n

Let M be the fourth test score.

Overall average \bar{x} = ( 60 + 78 + 86 + M) / 4

79.5 = ( 60 + 78 + 86 + M) / 4

Solve for M

79.5 = (224 + M) / 4

M = 94

Forth test score = 94